1. Special Venture Co. 10-year convertible bond pays a 6% semi-annual coupon (face value is $1000). The yield to maturity is 10% on debt of this quality. Surrendering the bond under the conversion terms allows the bondholder to receive 50 shares of stock. What is the market value of the convertible bond if the stock trades at $21 per share?

A. $750.76

B. $754.22

C. $1,050.00

D. $1,500.00

E. $1,754.22

2. Hartford Thatching is an unlevered firm with 30 million shares outstanding. The firm shares trade for $10 and the firm has a marginal tax rate of 40%. If the firm just was approved for a $50 million loan, then what is the firm worth after taking the loan?

A. $300 million

B. $320 million

C. $420 million

D. $470 million
Answer #1

1

Value of bond =max(price if converted,price of straight bond)

Price if converted = shares*price = 50*21 = 1050

Price of straight bond

                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =10x2
Bond Price =∑ [(6*1000/200)/(1 + 10/200)^k]     +   1000/(1 + 10/200)^10x2
                   k=1
Bond Price = 750.76

there fore value of bond = 1050

2

Value of levered firm = Value of unlevered firm + debt*tax rate = 300+50*0.4 = 320m